# High-frequency tails in spectral densities

### v1

@@ II.3 Model system
We use all five functional forms presented in Fig. 1 to include the eight modes (peaks) of the spectral density, resulting in five similar SDs (see Fig. 2). The solvent is included via DL functional form (i.e., setting $\omega_{k}=0$ in Eq. (11)) in each of the five spectral densities. Because the low-frequency range of the spectral density is dominated by this DL solvent feature, this ensures that it is virtually identical for the five SDs we are testing. The SD peaks appear between $\sim 300$ cm-1 and $\sim 4000$ cm-1. Here the SD’s obtained with single-peak functional forms (displaced Drude, the UBO and the RLC circuit) display nearly perfect agreement in the vicinity of each peak, while deviating slightly more at the peak edges. The tallest peaks of the SD extend to $J(\omega_{2})\approx 13,000$ and $J(\omega_{1})\approx 18,000$ respectively; we chose not to show the full $y$-range of the SD’s, as this would make the small differences between different functional forms shown on Fig. 2 indistinguishable. The two-peak functions (both Ohmic and Superohmic) visually display more pronounced deviations from the other three, but still match the peak positions, widths and intensities. The high-frequency range of the spectral density is shown on the inset and displays large relative differences between the SD’s constructed with RLC circuit, the displaced Drude and the UBO peaks, as we have seen on panel 1(e). Note that UBO and both of the two-peak functions yield virtually identical high-frequency tails, since all three functional forms decay faster than the DL solvent peak.
@@
The simulation is initialized in the product state $\rho(0)=\rho_{s}(0)\otimes\rho_{b}^{\beta}$, where $\rho_{s}(0)=\frac{(\ket{0}+\ket{1})(\bra{0}+\bra{1})}{2}$ is the reduced density matrix of the system at initial time and the bath is initially in thermal state at $300$ K. We integrate the master equation obtained based on time-dependent variational principle using the 4th order Runge-Kutta integrator with the 5th order error estimator, also known as the Dormand-Prince algorithm.[56] The timestep is set by the absolute and relative tolerance bounds of $10^{-8}$ and $10^{-5}$ respectively. We use the hierarchy cutoff of $25$ for each HEOM term, and add $8$ low temperature correction Padé terms.[57]
III Results and Discussion
@@ III.1 Dephasing dynamics
Figure 3: Dynamics of the model system in water using different spectral density basis functions. (a) Dephasing is insensitive to the choice of the spectral density basis functions. (b) Population relaxation is dominated by the high-frequency tails of the spectral density (SD). The color-coded rates of decay of SD peaks are shown for the different functional forms. Number 2 in the numerator of the RLC circuit scaling is to emphasize that it has a prefactor that’s double that of the displaced Drude oscillators (see Eq. (19)). The inset shows that even at early times (after $\sim 25$ fs) the population dynamics deviates significantly depending on the choice of the functional form of SD peaks.
@@
We first focus on the dephasing in the system, shown in Fig. 3a, where the absolute value of the off-diagonal element of the reduced density matrix of the system is plotted as a function of time with different functional forms of the SD peaks. The solvent alone (solid black curve in Fig. 3a) sets dephasing at $\sim 30$ fs. Inclusion of the eight vibrational modes in addition to the DL solvent feature increases the initial dephasing rate, but also introduces additional structure to the dephasing dynamics. More precisely, addition of sharp peaks in the range of frequencies between $\sim 300$ cm-1 and $\sim 2000$ cm-1 results in recurrences at $\sim 20$ fs and $\sim 35$ fs, that are not present with solvent-only DL bath. All of these observations echo the findings presented in Ref. 47.
@@
More significantly for this study, the dephasing is insensitive to the choice of the functional form of the vibrational modes in the SD, at least for the five functional forms we tested. The dephasing dynamics up to $\sim 20$ fs is determined by the peak positions, widths, and intensities[47] - all of which are the same for all the functional forms tested. At later times the decay of coherence is determined by the low-frequency end of the spectral density. Even though different functional forms have different low-frequency behaviors (see Eqs. (12), (15), (18) and (21)), their difference in contribution to the low-frequency end of the SD is masked by the DL solvent feature, which dominates in this frequency range (see Fig. 2).
@@
To summarize, we find that dephasing is insensitive to the precise functional form of the structured vibrational peaks in condensed phase dynamics both at early times (controlled by molecular vibrations) and at later times (controlled by the solvent).
III.2 Population relaxation
@@
Population relaxation (Fig. 3b) is about 3 orders of magnitude slower than dephasing for our model system. This corresponds to a typical nonradiative relaxation time of $1$ to $100$ ps for molecules in solution in the absence of near-resonant transition pathway, for example due to the presence of conical intersection. The solvent alone causes excited state population to decay with a lifetime (inverse of exponential decay rate) of $10.5$ ps. We observe that the rate of population relaxation is in some cases increased significantly by the addition of sharp peaks in the range of frequencies between $\sim 300$ cm-1 and $\sim 2000$ cm-1. The key finding of this study is that this effect differs significantly depending on the functional form chosen to represent the peaks. Addition of the RLC circuit peaks yields the fastest population decay with lifetime of $5.4$ ps, followed by the displaced Drude oscillator peaks with the lifetime of $7.0$ ps. Brownian and 2-peak Ohmic functions yield similar population lifetimes of $9.8$ ps, just a little faster than that caused by only the DL solvent feature.
@@
To explain these differences we refer back to Fig. 1. The different functional forms to represent each peak in the SD have identical reorganization energies, peak widths and characteristic frequencies, resulting in tiny absolute deviations in the vicinity of the peaks (panels a,b). Yet, these small differences in the functional form of the peaks affect population relaxation rates significantly. We interpret this to be caused by the large relative differences in the high-frequency tails of the five functional forms. The population relaxation rate within Born-Markov approximation[58] is primarily determined (Eq. (24)) by the SD value in the vicinity of the frequency of electronic transition $J(\Omega)$. For the model system we study $\Omega=35650$ cm-1. The Brownian and 2-peak Ohmic as well as superohmic tails decay quickly (as $\omega^{-3}$, $\omega^{-7}$ and $\omega^{-5}$ respectively, see Eqs. (16) and (22)). For these functional form choices the decay rate at the high-frequency end of the spectrum is higher than that of the DL solvent feature, which decays as $\omega^{-1}$, see Eq. (13). Therefore, the presence of these peaks does not significantly influence the high-frequency tails of the overall SD (see the right inset of Fig. 2), yielding the population lifetime that is only marginally smaller than the $10.5$ ps mark dictated by the solvent alone. This result is expected from mathematical considerations, but counter-intuitive and perhaps that’s why unappreciated in quantum dynamics community. To reiterate, the DL functional form used ubiquitously to represent the low-frequency solvent features in the spectral density not only determines the overall timescale for dephasing, but also has a dominant effect on the overall rate of population relaxation over molecular vibrations when they are represented in the SD via UBO functional form, as is customary.
@@
In contrast to UBO, both the displaced Drude oscillator and the RLC circuit functional forms have peak tails that decay as $\omega^{-1}$, the same rate as the solvent feature. Therefore, vibrations represented with these functional forms can have significant contribution to the overall SD at the high-frequency end of the spectrum. The relative importance of the solvent vs vibrations represented with either Drude or RLC functional forms depends also on the frequency of electronic transition, positions of vibrational peaks, and the difference between both reorganization energies and peak widths of the solvent vs vibrations. Note that the RLC circuits yield faster decay rate as compared to the displaced Drude oscillators because the prefactor of the dominant ($\omega^{-1}$) term at high frequency is two times larger for the former (see the inset of Fig. 2 and Eqs. (13), (19)).
@@
Thus, accurate calculations of slow electronic population relaxation require the precise functional form of the vibrational peaks in addition to the standard peak parameters (widths, positions and intensities). The latter can be obtained experimentally[47, 45, 46] or computationally,[44, 42, 43] while extracting the former from noisy data appears unfeasible at present. Moreover, even if the precise functional form is known, its representation within the numerically exact simulation can be challenging. For example, discretization of the SD or a high-frequency cutoff can introduce error from mishandling the high-frequency end of the SD. Additionally, efficiency considerations within HEOM calculations limit the choice of functional forms of the peaks. Finally, quantum analog simulator devices rely on physical processes to construct the bath, which set the functional form of the simulatable SD peaks. We suggest a recipe to account for the difference between the target and the simulatable SD functional forms in the following section.
III.3 Accounting for different SD tails
Absence of high-frequency (>$10,000$ cm-1) features in the SD ensures a separation of timescales by 1-3 orders of magnitude between fast dephasing and slow population relaxation. This timescale separation means that the two processes are (essentially) independent. In this limit the differences in population relaxation rates due to the different functional forms in the SD peaks can be accounted for using Eq. (23).
$p_{\text{target}}(t)\approx p_{eq}+\left(p_{\text{source}}(t)-p_{eq}\right)e^{-(\gamma_{\text{target}}-\gamma_{\text{source}})t}.$ (23)
Here $p(t)$ is the population of excited state as a function of time of the (known) source and the (unknown) target, which differ only by the functional form of the SD peaks, $p_{eq}=p_{\text{source}}(\infty)=p_{\text{target}}(\infty)$ is the population at thermal equilibrium, and $\gamma$ is the overall population decay rate,
@@ III.3 Accounting for different SD tails
obtained for both source and target within Born-Markov approximation (Eq. (24)).[58] This rate only depends on temperature, the frequency of electronic transition $\Omega$ and the value of the spectral density at that frequency. Eq. (23) is strictly true only when the source and the target only differ by the overall population decay rate and this difference grows exponentially with time. However, we find it to be an accurate approximation as long as the population relaxation is much slower than dephasing. We derived Eq. (23) for population relaxation obtained via Lindblad equation (see Appendix B), which predicts a simple exponential decay of initial condition to approach thermal equilibrium. For our model system the Lindblad prediction fails at short times, but accurately captures the long-time dynamics and approaches the exact (thermal equilibrium) result $p_{eq}$ as time tends to infinity. This suggests that we can use Eq. (23) together with population relaxation rates calculated based on Eq. (24) to accurately approximate the long-time trends in population relaxation dynamics for any known SD.
This is a useful result in the context of the recently proposed quantum analog simulator device,[38] which uses gate defined double quantum dots and the series of RLC circuits to simulate the system and bath parts of the spin-boson problem. As noted by the authors, the use of RLC circuits to represent the SD of the bath yields functional form of Eq. (17), which is similar (but not identical) to the UBO functional form (Eq. (14)). The ability to establish a simple connection between the functional form obtained from RLC circuits in quantum analog simulators and other target functional forms of the SD peaks is required to simulate systems, whose spectral density peaks do not decay as $\omega^{-1}$.
We illustrate this by taking the result of the HEOM simulation with RLC circuits bath and adjusting it using Eqs. (23) and (24). Fig. 4 shows the RLC circuit results adjusted to reproduce the displaced Drude and UBO peaks as black dotted lines. Recall that the population dynamics with RLC circuit peaks differs substantially from the other two (see Fig. 3); upon adjustment via Eq. (23) we recover excellent agreement with both target HEOM results. Note that the early-time accuracy of the adjusted result does not suffer from the inability of the Lindblad equation (used to derive Eq. (23)) to capture early-time oscillations in excited state population. While early-time dynamics of both target and source is not purely exponential, the oscillations are accurately captured by either of the functional forms, leaving only the overall population decay rate to be adjusted by Eq. (23).
@@
IV Conclusions and Implications

### v2

@@ 2.3 Model system
We use all five functional forms presented in Fig. 1 to include the eight modes (peaks) of the spectral density, resulting in five similar SDs (see Fig. 2). The solvent is included via DL functional form [that is, setting $\omega_{k}=0$ in Eq. (11)] in each of the five spectral densities. Because the low-frequency range of the spectral density is dominated by this DL solvent feature, this ensures that it is virtually identical for the five SDs we are testing. The SD peaks appear between $\sim 300$ cm-1 and $\sim 4000$ cm-1. Here the SD’s obtained with single-peak functional forms (displaced Drude, the UBO and the RLC circuit) display nearly perfect agreement in the vicinity of each peak, while deviating slightly more at the peak edges. The tallest peaks of the SD extend to $J(\omega_{2})\approx 13,000$ and $J(\omega_{1})\approx 18,000$ respectively; we chose not to show the full $y$-range of the SD’s, as this would make the small differences between different functional forms shown on Fig. 2 indistinguishable. The two-peak functions (both Ohmic and Superohmic) visually display more pronounced deviations from the other three, but still match the peak positions, widths and intensities. The high-frequency range of the spectral density is shown in the inset and displays large relative differences between the SD’s constructed with RLC circuit, the displaced Drude and the UBO peaks, as we have seen on panel 1(e). Note that UBO and both of the two-peak functions yield virtually identical high-frequency tails since all three functional forms decay faster than the DL solvent peak.
@@
The simulation is initialized in the product state $\rho(0)=\rho_{s}(0)\otimes\rho_{b}^{\beta}$, where $\rho_{s}(0)=\frac{(\ket{0}+\ket{1})(\bra{0}+\bra{1})}{2}$ is the reduced density matrix of the system at the initial time and the bath is initially in the thermal state at $300$ K. We integrate the master equation obtained based on the time-dependent variational principle using the 4th order Runge-Kutta integrator with the 5th order error estimator, also known as the Dormand-Prince algorithm.58 The timestep is set by the absolute and relative tolerance bounds of $10^{-8}$ and $10^{-5}$ respectively. We use the hierarchy cutoff of $25$ for each HEOM term, and add $8$ low temperature correction Padé terms.59
3 Results and Discussion
@@ 3.1 Dephasing dynamics
Figure 3: Dynamics of the model system in water using different spectral density basis functions. (a) Dephasing is insensitive to the choice of the spectral density basis functions. (b) Population relaxation is dominated by the high-frequency tails of the spectral density (SD). The color-coded rates of decay of SD peaks are shown for the different functional forms. Number 2 in the numerator of the RLC circuit scaling is to emphasize that it has a prefactor that’s double that of the displaced Drude oscillators [see Eq. (19)]. The inset shows that even at early times (after $\sim 25$ fs) the population dynamics deviates significantly depending on the choice of the functional form of SD peaks.
@@
We first focus on the dephasing in the system, shown in Fig. 3a, where the absolute value of the off-diagonal element of the reduced density matrix of the system is plotted as a function of time with different functional forms of the SD peaks. The solvent alone (solid black curve in Fig. 3a) sets dephasing at $\sim 30$ fs. Inclusion of the eight vibrational modes in addition to the DL solvent feature increases the initial dephasing rate, but also introduces additional structure to the dephasing dynamics. More precisely, the addition of sharp peaks in the range of frequencies between $\sim 300$ cm-1 and $\sim 2000$ cm-1 results in recurrences at $\sim 20$ fs and $\sim 35$ fs, which are not present with the solvent-only DL bath. All of these observations echo the findings presented in Ref. 43.
@@
More significantly for this study, the dephasing is insensitive to the choice of the functional form of the vibrational modes in the SD, at least for the five functional forms we tested. The dephasing dynamics up to $\sim 20$ fs is determined by the peak positions, widths, and intensities43 – all of which are the same for all functional forms tested. At later times the decay of coherence is determined by the low-frequency end of the spectral density. This is because the narrow high-frequency SD features alone lead to significant coherence recurrences after the initial fast decay (Fig. 3B of Ref. 43) due to wavepacket evolution in alternative potential energy surfaces. In contrast, the broad low-frequency solvent feature provides an irreversible decay path for the coherence that suppresses these recurrences and dictates the overall coherence loss. Although different functional forms have different low-frequency behaviors [see Eqs. (12), (15), (18) and (21)], their difference in contribution to the low-frequency end of the SD is masked by the DL solvent feature, which dominates in this frequency range (see Fig. 2).
@@
In summary, we find that dephasing is insensitive to the precise functional form of the structured vibrational peaks in condensed phase dynamics both at early times (controlled by molecular vibrations) and at later times (controlled by the solvent).
3.2 Population relaxation
@@
Population relaxation (Fig. 3b) is about 3 orders of magnitude slower than dephasing for our model system. This corresponds to the nonradiative relaxation time of several picoseconds – significantly slower than dephasing, yet fast enough that the radiative mechanism (i.e., fluorescense) can be ignored.60 The solvent alone causes excited state population to decay with a lifetime (inverse of exponential decay rate) of $10.5$ ps. We observe that the rate of population relaxation is in some cases increased significantly by adding sharp peaks in the range of frequencies between $\sim 300$ cm-1 and $\sim 2000$ cm-1. The key finding of this study is that this effect differs significantly depending on the functional form chosen to represent the peaks. The addition of the RLC circuit peaks yields the fastest population decay with a lifetime of $5.4$  ps, followed by the displaced Drude oscillator peaks with a lifetime of $7.0$  ps. Brownian and 2-peak Ohmic functions yield similar population lifetimes of $9.8$ ps, just a little faster than that caused by only the DL solvent feature.
@@
To explain these differences, we refer back to Fig. 1. The different functional forms that represent each peak in the SD have identical reorganization energies, peak widths, and characteristic frequencies, resulting in tiny absolute deviations in the vicinity of the peaks (panels a,b). However, these small differences in the functional form of the peaks significantly affect population relaxation rates. We interpret this to be caused by the large relative differences in the high-frequency tails of the five functional forms. The population relaxation rate within the Born-Markov approximation61 is mainly determined [Eq. (24)] by the SD value in the vicinity of the frequency of electronic transition $J(\Omega)$, which is significantly higher than molecular vibrations. Thus, for the model system we study, representing a typical electronic transition, the resonant contribution to electronic relaxation comes from the high-frequency tails of the SD.
The Brownian and 2-peak Ohmic as well as superohmic tails decay rapidly (as $\omega^{-3}$, $\omega^{-7}$ and $\omega^{-5}$ respectively, see Eqs. (16) and (22)). For these functional form choices the decay rate at the high-frequency end of the spectrum is higher than that of the DL solvent feature, which decays as $\omega^{-1}$, see Eq. (13). Therefore, the presence of these peaks does not significantly influence the high-frequency tails of the overall SD (see the right inset of Fig. 2), yielding the population lifetime that is only marginally smaller than the $10.5$ ps mark dictated by the solvent alone. This result is expected from mathematical considerations, but counterintuitive, and perhaps that is why unappreciated by the quantum dynamics community. To reiterate, the DL functional form used ubiquitously to represent the low-frequency solvent features in the spectral density not only determines the overall timescale for dephasing, but also has a dominant effect on the overall rate of population relaxation over molecular vibrations when they are represented in the SD via the UBO functional form, as is customary.
@@
In contrast to the UBO, both the displaced Drude oscillator and the RLC circuit functional forms have peak tails that decay as $\omega^{-1}$, the same rate as the solvent feature. Therefore, vibrations represented with these functional forms can have a significant contribution to the overall SD at the high-frequency end of the spectrum. The relative importance of the solvent vs vibrations represented with either Drude or RLC functional forms depends also on the frequency of electronic transition, positions of vibrational peaks, and the difference between both reorganization energies and peak widths of the solvent vs vibrations. Note that the RLC circuits yield a faster decay rate compared to the displaced Drude oscillators because the prefactor of the dominant ($\omega^{-1}$) term at high frequency is two times larger for the former [see the inset of Fig. 2 and Eqs. (13), (19)].
@@
Thus, for the population relaxation calculation presented here to be accurate, the precise functional form of the vibrational peaks is needed in addition to the standard peak parameters (widths, positions, and intensities). The peak parameters can be obtained experimentally43, 42, 40 or computationally,39, 37, 38, 41 but extracting the precise functional form from noisy data appears unfeasible at present. Moreover, even if the functional form of $J(\omega)$ is known, its representation within the numerically exact simulation can be challenging. For example, discretization of the SD or a high-frequency cut-off can introduce error from mishandling the high-frequency end of the SD. Additionally, efficiency considerations within the HEOM calculations limit the choice of functional forms of the peaks. Finally, quantum analog simulator devices rely on quantum hardware to construct the bath, which sets the functional form of the simulatable SD peaks. In the following section, we suggest a recipe to account for the difference between the real and the simulatable SD functional forms.
3.3 Correcting for the difference in high-frequency tails of the SD
We consider a situation in which true open quantum dynamics cannot be simulated due to physical or numerical constraints on the simulation, as is the case for HEOM, MCTDH, quantum analog simulation, and many other numerically exact methods. Instead, we have access to the simulated dynamics, which differs from the true dynamics only by the functional form of the SD peaks. Therefore, our goal is to recover the true population dynamics from the simulation results.
The most straightforward way to account for the difference in population relaxation rate is to add more flexibility to the fit by increasing the number of fit functions, that is, to add more features to the simulated SD. Since the functional form of any added feature in the simulated SD does not match that of the true SD, we do not expect this "brute force" strategy to yield a significant improvement in either the overall quality of the fit or in the resulting population relaxation dynamics.
Another potential approach would be to perform a weighted fitting of the real SD, such that the simulated SD is more accurately fit for some frequency ranges at the cost of larger errors elsewhere. For example, in the analog simulation given in Ref. 62 only near-resonance regions of the SD are fitted. This is justified since the effect of the environment on the dynamics of the system is not expected to be uniform across the frequency range. Such a fit could yield a better agreement in population relaxation dynamics, but it can be tricky to identify the optimal fitting weights for more complex dynamics.
Instead, we propose a simple, physically motivated, and universal approach, which can be used to adjust the results of the simulation after it has been completed. Our key assumption is that the dephasing dynamics is faster than the population relaxation dynamics by at least an order of magnitude, ensuring a separation of timescales between the two. We return with the justification of this assumption in models of realistic systems shortly. For now, we remark that this timescale separation means that the two processes are (essentially) independent. In this limit, differences in population relaxation rates due to the different functional forms in the SD peaks can be accounted for using Eq. (23).
$p_{\text{real}}(t)\approx p_{eq}+\left(p_{\text{simulated}}(t)-p_{eq}\right)e^{-(\gamma_{\text{real}}-\gamma_{\text{simulated}})t}.$ (23)
Here $p(t)$ is the population of excited state as a function of time with real and simulated SD; $p_{eq}=p_{\text{real}}(\infty)=p_{\text{simulated}}(\infty)$ is the population at thermal equilibrium, and
@@ 3.3 Correcting for the difference in high-frequency tails of the SD
is the overall population decay rate, obtained for both real and simulated SD within Born-Markov approximation [Eq. (24)].61 This rate only depends on temperature, the frequency of electronic transition $\Omega$ and the value of the spectral density at that frequency. Eq. (23) is strictly true only when the real and simulated SD’s only differ by the overall population decay rate and this difference grows exponentially with time. However, we find it to be an accurate approximation as long as the population relaxation is much slower than dephasing. We derived Eq. (23) for population relaxation obtained via Lindblad equation (see the SI), which predicts a simple exponential decay of initial condition to approach thermal equilibrium. For our model system the Lindblad prediction fails at short times, but accurately captures the long-time dynamics and approaches the exact (thermal equilibrium) result $p_{eq}$ as time tends to infinity. This suggests that we can use Eq. (23) together with population relaxation rates calculated based on Eq. (24) to accurately approximate the long-time trends in population relaxation dynamics for any known SD.
This is a useful result in the context of the recently proposed quantum analog simulator device,34 which uses gate-defined double quantum dots and the series of RLC circuits to simulate the system and bath parts of the spin-boson problem. As noted by the authors, the use of RLC circuits to represent the SD of the bath yields the functional form of Eq. (17), which is similar (but not identical) to the UBO functional form [Eq. (14)]. The ability to establish a simple connection between the functional form obtained from RLC circuits in quantum analog simulators and other target functional forms of the SD peaks is required to simulate systems, whose spectral density peaks do not decay as $\omega^{-1}$.
We illustrate this by taking the result of the HEOM simulation with the RLC circuit bath and adjusting it using Eqs. (23) and (24). Figure 4 shows the RLC circuit results adjusted to reproduce the displaced Drude and UBO peaks as black dotted lines. The population dynamics with RLC circuit peaks differs substantially from the other two. However, upon adjustment through Eq. (23) we recover excellent agreement with both HEOM target results. Note that the early-time accuracy of the adjusted result does not suffer from the inability of the Lindblad equation [used to derive Eq. (23)] to capture early-time oscillations in the excited-state population. This is because the oscillations in the early-time dynamics are accurately captured by either of the functional forms, leaving only the overall population decay rate to be adjusted by Eq. (23).
@@
3.4 Implications and Limitations of our Findings
In Sec. 3.2, we showed and rationalized the importance of the high-frequency tails in the SD for electronic population relaxation in an instructive model system. Based on this analysis, we concluded that the high-frequency tails will dominate the nonradiative relaxation when the corresponding spectral density has no features in the vicinity of the system transition frequency. That is, when
$\max_{k}(\omega_{k}+\gamma_{k})\ll\Omega.$ (25)
In this regime, the precise functional form of the SD becomes important to correctly predict the relaxation dynamics. The natural question is: when do these findings become relevant to predict the relaxation behavior of molecules?
We limit our considerations to electronic transitions that preserve spin symmetry, such that there are only two competing relaxation mechanisms – the radiative (fluorescense) and the nonradiative (internal conversion). Our analysis does not impact the description of radiative relaxation, as the spectral density for radiation fields $J(\omega)\propto\omega^{3}$ does not feature a cutoff frequency, so condition Eq. (25) cannot be satisfied.
For the non-radiative mechanisms, our analysis pertains to multiphonon electronic relaxation processes in molecules. These processes proceed via phonon-mediated near-resonant transitions from the low vibrational energy levels of the excited electronic state to multiple highly excited vibrational levels of the ground electronic state. The excess vibrational energy subsequently relaxes due to rapid intramolecular vibrational energy redistribution and collisions with the solvent60, 63, 64, 65 These processes can be effectively modeled as a two-level system bilinearly coupled to a bosonic bath.60 That is, they are encoded in the bath SD and fall within the scope of our analysis.
Our considerations do not apply to electronic relaxation promoted by strong coupling to anharmonic degrees of freedom of the bath, for instance, via conical intersections. In such cases, alternative computational approaches that can handle non-Gaussian environments are required such as the Multi-Configurational Time Dependent Hartree,4 semiclassical methods66 and the Automated Compression of Environments method.67
When constructing the model described in section 2.3 we also made several simplifying assumptions, which we do not expect to affect the validity of our conclusions. First, we truncated the electronic space to two states. Higher excited states can be added as needed, in which case the condition described by Eq. (25) must be fulfilled for the dominant relaxation path. Additionally, to reduce the numerical cost, we used the spectral density extracted from pure dephasing experiments to induce both dephasing and relaxation by coupling the system to a single bath of harmonic oscillators. We do not expect the split into two independent baths to change our conclusions, since the effects of the $\sigma_{x}$ and $\sigma_{z}$ couplings for the set of parameters we have considered are essentially decoupled via separation of timescales. Furthermore, while the dissipative SD is expected to be different from that responsible for pure dephasing, we expect that it will also satisfy Eq. (25) since it originates from the same environment.
To summarise, we expect our finding to be applicable to real molecules with slow nonradiative relaxation. That is, if: (i) the spin-boson model can be used to describe the process; and (ii) electronic and vibrational energy scales are separated as described in Eq. (25). In practice, these conditions are fulfilled by many molecules that relax nonradiatively with timescales of picosecond or longer.60 In these cases, Eq. (23) provides a convenient way to compare simulations of the same system obtained by using different spectral density functions.
4 Conclusions
The precise structure of the high-frequency tails of the spectral density (SD) peaks is typically considered to be inconsequential for open quantum dynamics. In contrast, here we show that different peak shapes lead to identical dephasing dynamics but different relaxation rates in a spin-boson calculation when the transition frequency of the system is significantly larger than the highest frequency of the bath. These discrepancies arise because in this regime the population relaxation is controlled by the high-frequency tails, which are several orders of magnitude smaller than the main SD features. In such cases the full decoherence dynamics requires an accurate representation of the high-frequency tails, in addition to the main features and the low-frequency behavior of the spectral density.